When the power is turned off on a turntable spinning at 65.0 rpm , you find that it takes 11.3 revolutions for it to stop while slowing down at a uniform rate.

What is the angular acceleration (in rad/s2), of this turntable?

How long does it take to stop after the power is turned off?
Answer #1

11.3 revolution = 11.3*2pi rad = 71 rad

N= 65 rpm

angular speed = 2pi*N / 60

= 2pi*65/60

= 6.8 rad/s

using the eqaution of angulat motion we have

w^2 - wo^2 = 2*

0 - 6.8^2 = 2**71

= 0.326 rad/s2

w= wo + t

t = (w - wo) /

= ( 0 - 6.8) / -0.326

= 20.84 s